OpenStudy (anonymous):

x+2/6 = 2/x-1 pleaseeeee help me !

OpenStudy (anonymous):

Is this the problem: \[\frac{x + 2}{6} = \frac{2}{x - 1}\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Just cross multiply. \[(x + 2)(x - 1) = 2(6)\] \[(x + 2)(x - 1) = 12\] \[x^{2} + x - 2 = 12\] \[x^{2} + x - 14 = 0\] Can you solve the quadratic now?

OpenStudy (anonymous):

no i forgot how to do that

OpenStudy (anonymous):

Just solve using the quadratic formula: \[x = \frac{-b ± \sqrt{b^{2} - 4ac}}{2a}\] Where: a = 1 b = 1 c = -14

OpenStudy (anonymous):

i learned it an easier way ,

OpenStudy (anonymous):

what 2 #'s multpliy to 1 and add up to 14?

OpenStudy (anonymous):

That's why we can't do that. There aren't any numbers that can work like that.

OpenStudy (anonymous):

ohhhh got it, well we haven't learned that formula so i will leave it like that

OpenStudy (anonymous):

Okay, complete the square? x^2+x−14=0 x^2+x=14 x^2+x+0.25=14+0.25 x+0.5=±3.77 Can you solve from here?

OpenStudy (anonymous):

Arte you sure you haven't learned any of this stuff?